### v1

@@ Abstract
Streams of data have become the ubiquitous communication model on today’s Internet. For strong anonymous communication, this conflicts with the traditional notion of single, independent messages, as assumed e.g. by many mixnet designs. In this work, we investigate the anonymity factors that are inherent to stream communication. We introduce progressive pruning, a methodology suitable for estimating the anonymity level of streams. By mimicking an intersection attack, it captures the susceptibility of streams against traffic analysis attacks. We apply it to simulations of tailored examples of stream communication as well as to large-scale simulations of Tor using our novel TorFS simulator, finding that the stream length, the number of users, and how streams are distributed over the network have interdependent impacts on anonymity. Our work draws attention to challenges that need to be solved in order to provide strong anonymity for stream-based communication in the future.
@@
anonymous communications, stream communication, anonymity metrics, intersection attacks, Tor, mixnets
@@ 1 Introduction
Anonymous communication is becoming increasingly important in today’s society, aiming to break sender-receiver linkability between communication partners. Like other types of online communication, it can either be carried out in the form of independent messages, or as streams, i.e. sequences of consecutive messages exchanged between the same sender-receiver pair. Today, data streams constitute the prevalent communication model underlying most practical applications on the Internet. In contrast, the traditional approach of anonymity systems regards independent messages as the base unit of communication. Attempts to break the users’ anonymity are typically carried out by a global or partially global adversaries as traffic analysis attacks. In such attacks, the adversary monitors the (encrypted) traffic and, based on the observed timing information, tries to correlate and thus link senders and receivers.
In this paper, we establish a more thorough understanding of how stream-based communication influences susceptibility to traffic analysis attacks. For this, we introduce progressive pruning, a methodology for estimating the anonymity level against intersection attacks, tailored to analyzing streams of communication. It works by combining the individual anonymity sets of messages as they occur over the progress of a stream, into a common pruned anonymity set for the whole stream. We argue that the perspective of progressive pruning is important in order to design future anonymity systems such that they can provide strong anonymity for stream communication. Progressive pruning enables estimation of the success chances of an intersection attack on the stream monitored by an attacker. For example, it can be used to compare the susceptibility to intersection attacks between different system variants.
We make use of this approach to derive general, asymptotic anonymity trends from stream properties. Our results indicate that anonymity is primarily influenced by the relative stream length among all streams, as well as the number of streams. We show that anonymity is distributed very unevenly and long streams implicitly take the role of cover traffic for others while experiencing poor anonymity themselves.
@@
  • We develop a methodology called progressive pruning (Section 3) to systematically evaluate the anonymity properties of streams, i.e. sequences of messages, as observed by an adversary who monitors traffic entering and leaving an ACN. We also discuss anonymity properties that can be derived.
@@
2 Preliminaries
2.1 Background and Problem Statement
Figure 1: Anonymity of stream-based communication. For some user (yellow), an attacker would like to determine the communication partner (red or blue). A typical approach for de-anonymization would be an intersection attack.
Anonymous communication has been an active field of research for many years, starting from the original Chaumian mix design [chaum1981untraceable] and branching into various directions that aim to establish anonymity in online communication by following different strategies [edman2009anon]. One aspect that is of particular significance in this regard, but has received little attention, is the communication model employed by the different approaches. In general, these can be categorized into two distinct classes: On the one hand, the traditional approach assumes the notion of single independent messages that are transferred. This aligns well with the long time prevalent asynchronous communication based e.g. on email. On the other hand, anonymous communication can be regarded as streams of communication between senders and receivers as depicted in Figure 1. This understanding allows for synchronous, interactive communication and aligns more closely with modern online communication habits such as web browsing. These two different strategies can also be seen when looking at specific anonymous communication networks (ACNs). While mixnets such as Loopix [piotrowska2017loopix] generally transfer single messages over independent paths through the network, onion routing systems like Tor [dingledine2004tor] carry streams of messages (i.e. a stream of bytes, divided into equally-sized cells in the case of Tor). The two communication models—independent messages vs. stream-based communication—constitute fundamentally incompatible, orthogonal principles. This becomes apparent both when looking at the (performance) costs per message and especially when considering the security and anonymity implications. Mixnets like Loopix rely heavily on the fact that multiple messages between the same sender-receiver pair take independent paths through the network [piotrowska2017loopix].
Yet, we observe that there is a lack of research investigating the inherent factors that influence anonymity in stream-based anonymous communication. Many publications (implicitly) rely on the independence of transferred messages to derive anonymity properties—in particular, the unlinkability of multiple messages between the same sender-receiver pair. At the same time, when analyzing specific anonymity systems for interactive communication, shortcomings in the achieved anonymity are easily attributed to the low-latency nature of the system and lack of mixing, resulting in in the susceptibility against traffic analysis attacks. In contrast, we argue that it is crucial to understand the inherent anonymity implications of stream-based communication itself, independently of other system properties such as latency or cover traffic. This is because stream-based communication is at the core of modern online communication. Our work therefore aims to shed light on basic, inherent anonymity properties of stream-based communication that are relevant not only for onion routing systems but also to other ACNs aiming to support stream-based communication. One step towards understanding such relations is to measure or estimate the achieved anonymity. While a plethora of anonymity metrics exists for many different applications [DBLP:journals/csur/WagnerE18], these are not primarily useful for understanding the (asymptotic) implications of stream-based communication.
Our goal therefore is to quantify the anonymity level of streams in a way that helps understand inherent factors which increase or reduce anonymity.
2.2 Adversary Model
Discussing and quantifying anonymity only makes sense with regard to a concrete adversary model. In order to evaluate the anonymity properties of stream-based communication, we assume that an attacker can monitor both ends of the streams. That is, the attacker can observe when packets of each stream are sent and received. This setting therefore focuses on traffic analysis attacks in which the attacker compares temporal traffic patterns at both ends of the streams, trying to identify pairs of communication. In the context of stream-based communication the primary attack is an intersection attack [berthold2001disadvantages], which will therefore be the main concern of this work. The idea behind such attacks is to exploit the knowledge of multiple messages belonging together by intersecting their anonymity sets in order to identify the communication partners. Our adversary model is inspired by a global passive adversary, which appears to be a sensible assumption today [nsa-buys-data], and deliberately abstracts from any concrete underlying communication system to capture the anonymity properties inherent to streams.
2.3 Related Work
The long research history of anonymous communication begins in 1981, when Chaum introduced the concept of a mix [chaum1981untraceable]. Even in this early work, the central idea to achieve unlinkability revolves around the notion of independent messages that are uniformly sized and thus indistinguishable for an outside observer. This assumption has since become a core principle of mixnet designs, up to modern mixnet systems like Loopix [piotrowska2017loopix] and Nym [diaz2021nym].
The idea to anonymize streams of data was first followed in the context of ISDN telephony [pfitzmann1991isdn]. The system primarily exploited the fixed-rate nature of the ISDN network to achieve traffic analysis resistance. With the introduction of onion routing [syverson1997anonymous], the stream notion gained further attention, resulting in its adoption as a primary design principle in Tor [dingledine04tor]. Other low-latency anonymity networks have since continued to utilize this approach [chen2015hornet, DBLP:journals/popets/ChenP17].
However, streams of data being transferred in a low-latency fashion bring with them increased susceptibility to traffic analysis attacks, in particular intersection attacks [berthold2001disadvantages]. This is also the fundamental approach followed by our progressive pruning methodology, whereas our work aims specifically at making the impacts of specific stream properties comprehensible. A version of this attack that instead operates on the basis of probabilities has been put forward as the statistical disclosure attack. In general, the field of traffic analysis attacks has been the subject of many research studies. Correlation techniques that have been proposed range from simple approaches like packet counting [DBLP:conf/ih/BackMS01] through information-theoretic estimates like cross-correlation coefficients [levine2004timing] and mutual information [DBLP:conf/pet/ZhuFGBZ04] to advanced matching techniques based on deep learning [DBLP:conf/ccs/NasrBH18]. While these methods constitute relevant real-world attacks, they are not necessarily useful to understand the essence of anonymity within stream-based communication, as influenced e.g. by the length of streams or the number of simultaneous streams. We therefore propose an alternate methodology that supports establishing an understanding of how different inherent stream properties influence anonymity.
@@ 3.1 Overview
We now introduce progressive pruning, a new methodology to analyze the anonymity implications of streams of data, as they are transferred via an ACN. We understand a stream as a reliable, unidirectional sequence of messages that are exchanged between the same pair of sender and receiver, over the same communication channel (e.g. through an ACN). Streams do not necessarily need to be order-preserving, however, messages need to be delivered within some reasonably bounded time. This transfer of streams of data comes with an increased vulnerability against intersection attacks. In such attacks, an attacker exploits the fact that multiple messages belong to the same communication partners by intersecting the anonymity set to reduce the number of candidates, keeping only candidates that are in line with all observations. Since the vulnerability to intersection attacks is what fundamentally differentiates streams from single messages, we design our method similarly. This way, we can identify anonymity effects that primarily stem from the stream characteristics, compared to the exchange of a single message.
The high-level goal of progressive pruning is as follows: Given a network trace of stream messages along with observation timestamps at its source and destination, we want to calculate an anonymity estimate for each of the contained streams. For this, progressive pruning builds upon the well-known notion of anonymity sets. However, it adds a time component: Initially, there may be a broad set of candidates, so the first message may have a large anonymity set. However, as communication proceeds over more messages, the anonymity set shrinks as candidates drop out. This may be the case because at some point their observed behavior may not be compatible with other observations. The fact that the set of candidates is pruned progressively as the message streams advance is what gives our methodology its name.
It should be noted that, as a consequence of the design, our anonymity metric does not provide an absolute value of system security or anonymity, which would be very hard or likely impossible. However, it can be used to analyze the relative changes to anonymity induced by different system parameters.
3.2 Detailed Description
@@
We assume that the input stream traces stem from a communication system of the following abstract notion: Unidirectional streams of messages are exchanged between senders and receivers. Messages can only be linked based on timing information and share a uniform size. We further assume that there is no message loss, and the delay of each message (the time between being sent and received) is bounded between $[D_{min},D_{max}]$. Depending on whether we want to evaluate the anonymity level of senders or receivers, either (1) senders send all their messages to the same receiver, or (2) receivers receive all their messages from the same sender. The attacker’s goal is to link senders and receivers. For clarity, we stick to the sender perspective from now on. Consequently, the attacker wants to find the corresponding receiver of each sender. The opposite direction is fully analogous. As our deliberately simple attacker represents a concrete attack instance, the resulting anonymity values can be interpreted as upper bounds on the achieved anonymity. Our assumptions regarding the stream model are not meant to mimic any existing real-world system, but rather allow investigating the inherent properties of stream communication.
Figure 2: Anonymity set calculation for individual messages of a sending stream. For each message from the sender (yellow), the set of potentially corresponding receiver messages is determined based on the lower and upper network delay bounds $D_{min}$ and $D_{max}$. Receivers of messages within this window (red, blue) may be communication partners of the sender. The anonymity set of the whole stream is then calculated by intersecting the individual message anonymity sets.
@@
Let $S$ be the set of senders and $R$ the set of receivers. Accordingly, let $M_{S}$ be the set of messages sent by all senders, and $M_{R}$ the set of messages received by all receivers. $M_{S}[s]$ is the set of all messages sent by sender $s\in S$ and $M_{R}[r]$ is the set of all messages received by receiver $r\in R$. By $S[m]$, we denote the sender $s\in S$ of message $m$. Analogously, $R[m]$ is the receiver $r\in R$ of message $m$. The relationship between sent and received messages can be characterized as follows: $\forall m_{S}\in M_{S}:\exists m_{R}\in M_{R}:m_{S}=m_{R}$. It is the adversary’s goal to find this mapping, i.e. determine the receiver of a sent message, breaking sender-receiver unlinkabilty. Furthermore, let $T_{S}[m]$ be the time when message $m$ was sent, and $T_{R}[m]$ the time when message $m$ was received.
@@
$\Psi_{MM}[m_{S}]$ =$=\{m_{R}$ $\in M_{R}\mid T_{R}[m_{R}]\in[T_{S}[m_{S}]+D_{min}~{},~{}T_{S}[m_{S}]+D_{max}]\}$
Based on this, it is straightforward to define the message-receiver anonymity set $\Psi_{MR}$, of a message, denoting the set of potential receiver candidates, a subset of $R$:
@@
We now extend these anonymity set definitions to multiple messages from the same sender (a stream). By doing so, we want to capture the anonymity set of senders as a whole, taking all their messages into account. The simplest approach would be to simply intersect the message-receiver anonymity sets for all messages from a sender: $\bigcap_{m_{S}\in M_{S}[s]}\Psi_{MR}[m_{S}]$. However, this falls short of the fact that the adversary can trivially also exploit the number of messages to differentiate between candidates. For example, if a sender sent 10 messages in a short period of time, but a receiver received only one message that falls into the message-message anonymity sets of all sent messages, the simple intersection approach would not exclude that receiver although it obviously cannot be a candidate because it received too few messages.
We solve this issue by formulating the following assignment problem: Let $s$ be a sender of messages $M_{S}[s]$ and $r$ a receiver of messages $M_{R}[r]$. Then, $r$ can only be a candidate as the receiver of the messages from $s$, if there exists a mapping $\sigma:M_{S}[s]\to M_{R}[r]$ that assigns each sent message to a corresponding received one and meets the following two requirements: (1) If $\sigma(m_{S})=m_{R}$, then $m_{R}\in\Psi_{MM}[m_{S}]$ (the messages potentially correspond with respect to the delay constraints), and (2) $\sigma$ is injective (each received message is assigned at most once).
Given the sets of sent and received messages, it is possible to determine whether such a mapping exists, e.g. in a brute-force manner. Moreover, in Section 3.4 we will present an efficient algorithm to solve this task in linear time. If a mapping that meets the the aforementioned requirements can be found, we say that $r$ is a candidate for $s$. Consequently, we can define the sender-receiver anonymity set of sender $s$ as follows:
@@ 3.3 Model Discussion
Intuitively, we can understand the presented construction as the anonymity set of a sender $s$ over time. When taking into account only a single message, then all receivers that have received a message in the respective timeframe are possible communication partners of $s$. Additional messages in the stream, however, add additional constraints that can be exploited by the adversary. As such, more messages will likely result in a shrinking anonymity set (if the streams of communication are not synchronized). At the same time, the number of observed messages that are necessary to break anonymity can be regarded as an indicator for the costs an adversary has to invest for the attack.
Another observation that can be made is that the number of messages sent by a sender is highly relevant for the resulting anonymity set. In particular, if the streams differ in their number of messages, then the stream length is useful for an adversary to differentiate the senders. While this comes at no surprise, it has an interesting implication: A sender that sends more data than anyone else does not retain any anonymity at all. In general, the shorter the communication is in comparison to the others, the more likely it is that a sender can “hide” among the other senders. We formally verified this important relationship and present the analysis in the appendix. Additionally, the absolute latency value of messages does not directly make a difference. Instead, an attacker only benefits from a low deviation of latency (jitter) because it makes the traffic more predictable, allowing use of a smaller window.
3.4 Algorithm: Matching Candidate Messages

### v2

@@ Abstract
Stream-based communication dominates today’s Internet, posing unique challenges for anonymous communication networks (ACNs). Traditionally designed for independent messages, ACNs struggle to account for the inherent vulnerabilities of streams, such as susceptibility to intersection attacks. In this work, we address this gap and introduce progressive pruning, a novel methodology for quantifying the susceptibility to intersection attacks. Progressive pruning quantifies and monitors anonymity sets over time, providing an assessment of an adversary’s success in correlating senders and receivers. We leverage this methodology to analyze synthetic scenarios and large-scale simulations of the Tor network using our newly developed TorFS simulator. Our findings reveal that anonymity is significantly influenced by stream length, user population, and stream distribution across the network. These insights highlight critical design challenges for future ACNs seeking to safeguard stream-based communication against traffic analysis attacks.
@@
anonymous communications, intersection attacks, Tor
@@ 1 Introduction
Anonymous communication is becoming increasingly vital in today’s society, as it aims to prevent the correlation of senders and receivers in communication networks. While early anonymous communication networks (ACNs) focused on high-latency networks operating with independent messages, the dominant model today involves low-latency networks designed to transport streams of data. A notable example is the Tor network [dingledine04tor]. However, a fundamental vulnerability of such systems lies in traffic analysis attacks. In these attacks, adversaries monitor (encrypted) traffic and use observed timing information to correlate and link senders with receivers. Recent incidents, such as the German Federal Criminal Police Office (BKA) successfully executing a timing attack [panorama-boystown], highlight the pressing need to address these vulnerabilities. Dismissing such attacks as out of scope is no longer a viable option.
In this paper, we establish a deeper understanding of how stream-based communication impacts susceptibility to traffic analysis. For this, we introduce progressive pruning, a methodology for quantifying anonymity sets under intersection attacks, specifically designed for analyzing streams of communication. Progressive pruning works by combining anonymity sets of individual messages as they occur over the course of a stream, into a common pruned anonymity set for the entire stream. This approach enables estimating an attacker’s success chances when performing an intersection attack on a monitored stream. For instance, progressive pruning can be used to compare the susceptibility of different system variants to such attacks. We argue that adopting this perspective is crucial for designing future anonymity systems capable of providing robust anonymity for stream-based communication.
We make use of progressive pruning to derive general, asymptotic anonymity trends from stream properties. Our results indicate that anonymity is primarily influenced by the relative stream length among all streams, as well as the number of streams. We show that long streams implicitly take the role of cover traffic for others while experiencing poor anonymity themselves. Applying progressive pruning to Tor as a real-world network, we find that anonymity is distributed very unevenly.
@@
  • We develop a methodology called progressive pruning (Section 3) to systematically evaluate the anonymity properties of streams, i.e., sequences of messages, as observed by an adversary who monitors traffic entering and leaving an ACN. We also discuss anonymity properties that can be derived.
@@
2 Related Work
The study of anonymous communication began in 1981, when Chaum introduced the concept of a mix [chaum1981untraceable]. The key idea of achieving unlinkability through independent messages of uniform size, making them indistinguishable to observers, has since become foundational for mixnet designs, including modern systems like Loopix [piotrowska2017loopix] and Nym [diaz2021nym]. The anonymization of streams of data gained prominence with the introduction of onion routing [syverson1997anonymous], which became a core design principle in Tor [dingledine04tor]. Other low-latency anonymity networks have since adopted this approach [chen2015hornet, DBLP:journals/popets/ChenP17].
@@ 3 Progressive Pruning
In this section, we introduce progressive pruning, a novel methodology designed to analyze the anonymity implications of data streams in anonymous communication networks (ACNs).
3.1 Adversary Model
In order to evaluate the anonymity properties of stream-based communication, we assume that an attacker can monitor both ends of the streams. That is, the attacker can observe when packets of each stream are sent and received. This setting, therefore, focuses on traffic analysis attacks in which the attacker compares temporal traffic patterns at both ends of the streams, trying to identify pairs of communication (see Figure 1). In the context of stream-based communication, the primary attack is an intersection attack [berthold2001disadvantages], which will therefore be the main concern of this work. The idea behind such attacks is to exploit the knowledge of multiple messages belonging together by intersecting their anonymity sets in order to identify the communication partners. Our adversary model is inspired by a global passive adversary, which appears to be a sensible assumption today [nsa-buys-data], and deliberately abstracts from any concrete underlying communication system to capture the anonymity properties inherent to streams.
Figure 1: For some user (yellow), an attacker aims to identify the communication partner (red or blue) using traffic analysis.
@@ 3.2 Overview
Progressive pruning builds on the well-known notion of anonymity sets—but evolves over time: Given a network trace of messages with observation timestamps at their source and destination, we model the core concept of an intersection attack and quantify anonymity sets per stream. Initially, a message may have a large anonymity set as there may be a broad set of candidates. However, as communication progresses and more messages are exchanged, the anonymity set shrinks as candidates are excluded. For example, the observed timing of events may conflict with other previously gathered observations, rendering them incompatible. Pruning the set of candidates progressively gives our methodology its name. We illustrate progressive pruning for a single message in Figure 2.
We note that, as a consequence of the design, our anonymity metric does not provide an absolute value of system security or anonymity, which would be very hard or likely impossible. However, it can be used to analyze the relative changes to anonymity induced by different system parameters.
3.3 Modeling Intersection Attacks
@@
We assume that the input stream traces stem from a communication system of the following abstract notion: Unidirectional streams of messages are exchanged between senders and receivers. Messages can only be linked based on timing information and share a uniform size. We further assume that there is no message loss, and the delay of each message (the time between being sent and received) is bounded between $[D_{min},D_{max}]$. Depending on whether we want to evaluate the anonymity of senders or receivers, either (1) senders send all their messages to the same receiver, or (2) receivers receive all their messages from the same sender. The attacker’s goal is to link senders and receivers. For clarity and without loss of generality, we focus on the sender’s perspective. Consequently, the attacker aims to identify the corresponding receiver for each sender. As our deliberately simple attacker represents a concrete attack instance, the resulting anonymity can be interpreted as upper bounds on the achieved anonymity. Our assumptions for the stream model allow investigating the inherent properties of stream communication but are not intended to replicate any real-world system.
@@
Let $S$ be the set of senders and $R$ the set of receivers. Accordingly, let $M_{S}$ be the set of messages sent by all senders, and $M_{R}$ the set of messages received by all receivers. $M_{S}[s]$ is the set of all messages sent by sender $s\in S$ and $M_{R}[r]$ is the set of all messages received by receiver $r\in R$. By $S[m]$, we denote the sender $s\in S$ of message $m$. Analogously, $R[m]$ is the receiver $r\in R$ of message $m$. The relationship between sent and received messages can be characterized as follows: $\forall m_{S}\in M_{S}:\exists m_{R}\in M_{R}:m_{S}=m_{R}$. It is the adversary’s goal to find this mapping, i.e., determine the receiver of a sent message, breaking sender-receiver unlinkabilty. Furthermore, let $T_{S}[m]$ be the time when message $m$ was sent, and $T_{R}[m]$ the time when message $m$ was received.
@@
$\Psi_{MM}[m_{S}]=\{m_{R}\in M_{R}\mid T_{R}[m_{R}]\in[T_{S}[m_{S}]+D_{min}~{},~{}T_{S}[m_{S}]+D_{max}]\}$
We can use this to define the message-receiver anonymity set $\Psi_{MR}$ for a message, which represents the potential receiver candidates and forms a subset of $R$:
@@
Figure 2: Progressive Pruning: For each sender message (yellow), potential receiver messages are identified based on the network delay bounds $D_{min}$ and $D_{max}$. Receivers of messages within this window (red, blue) are considered as candidates. The stream’s anonymity set is determined by intersecting these message anonymity sets over time.
@@
We now extend these anonymity set definitions to multiple messages from the same sender—a stream. By doing so, we want to capture the anonymity set of senders as a whole, taking all their messages into account. The simplest approach would be to simply intersect the message-receiver anonymity sets for all messages from a sender: $\bigcap_{m_{S}\in M_{S}[s]}\Psi_{MR}[m_{S}]$. However, this falls short of the fact that the adversary can trivially also exploit the number of messages to differentiate between candidates. For example, if a sender sent 10 messages in a short period of time, but a receiver received only one message that falls into the message-message anonymity sets of all sent messages, the simple intersection approach would not exclude that receiver although it obviously cannot be a candidate because it received too few messages.
We address this issue by formulating the following assignment problem: Let $s$ be a sender of messages $M_{S}[s]$ and $r$ a receiver of messages $M_{R}[r]$. Then, $r$ can be a candidate receiver for the messages from $s$ only if there exists a mapping $\sigma:M_{S}[s]\to M_{R}[r]$ that assigns each sent message to a corresponding received one and meets the following two requirements: (1) If $\sigma(m_{S})=m_{R}$, then $m_{R}\in\Psi_{MM}[m_{S}]$ (messages potentially correspond w.r.t. the delay constraints), and (2) $\sigma$ is injective (each received message is assigned at most once).
Given the sets of sent and received messages, it is possible to determine whether such a mapping exists, e.g., in a brute-force manner. In Section 3.5, we will present an efficient algorithm to solve this task in linear time. If a mapping that meets the the aforementioned requirements can be found, we say that $r$ is a candidate for $s$. Consequently, we can define the sender-receiver anonymity set of sender $s$ as follows:
@@ 3.4 Model Discussion
Intuitively, we can understand the presented construction as the anonymity set of a sender $s$ over time. When taking into account only a single message, then all receivers that have received a message in the respective timeframe are possible communication partners of $s$. Additional messages in the stream, however, add additional constraints that can be exploited by the adversary. As such, more messages will eventually decrease the anonymity set. The number of observed messages that are necessary to break anonymity can be regarded as an indicator for the costs an adversary has to invest for the attack.
Another observation that can be made is that the number of messages sent by a sender is highly relevant for the resulting anonymity set. In particular, if the streams differ in their number of messages, then the stream length is useful for an adversary to differentiate senders. While this comes at no surprise, it has an interesting implication: A sender that sends more data than anyone else does not retain any anonymity. In general, the shorter the communication is in comparison to the others, the more likely it is that a sender can “hide” among the other senders. We also verified this important relationship formally [progressive-pruning-arxiv]. Additionally, the absolute latency value of messages does not directly make a difference. Instead, an attacker only benefits from a low deviation of latency (jitter) because it makes the traffic more predictable, allowing use of a smaller window.
3.5 Algorithm for Matching Candidate Messages